Two mathematics classes took the same test. The mean of the first class of twenty students was 90%. The mean of the second class was 75%. If the mean grade for the combined classes was 81%, how many students were in the second class?
(A) 20 (B) 30 (C) 23 (D) 22 (E) 28

Answer by richwmiller(17219)

f/20=90
s/y=75
(f+s)/(20+y)=81
f = 1800, s = 2250, y = 30
The second class had 30 students
(B)30